OpenStudy (anonymous):

my teacher want us to find the derivative of this problem using the limit rule, but I get lose using the limit rule: (1/(x+1))

OpenStudy (anonymous):

@helder_edwin

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

but when u take the derivative, u get -1/(x+1)^2. but thanks for the response

OpenStudy (anonymous):

Yeah, I have done mistake. This is the new answer. Sorry.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

I forgot to change \[\frac{ d }{ dx }\left( \frac{ 1 }{ x } \right)\] It should be \[\frac{ d }{ dx }\left( \frac{ 1 }{ x+1 } \right)\]